OpenStudy (thecuteone25):

will MEDAL and FAN anyone who helps me the most Prove that the decimal number 6.52 is rational by finding its fractional form. * 6 52/10 652/100 65 2/100 652/1,000

OpenStudy (thecuteone25):

@Hope210 @mathmale

OpenStudy (hope210):

@SolomonZelman

OpenStudy (thecuteone25):

@Alphabet_Sam

OpenStudy (anonymous):

To make a decimal into a fraction, you have to think about how many times it needs to be multiplied by 10 to stop being a decimal. I'll show you two examples: Let's take the number 4.1. In this case, it needs to be multiplied by 10 once, so: 4.1*10=41 Now take this integer that you got (in this case 4.1) and divide it by all the tens you needed multiplied together. So: 4.1=41/10 Another example with the number 7.6593 Multiply by 10 until it has no decimal: 7.6593*10=76.593 76.593*10=765.93 765.93*10=7659.3 7659.3*10=76593 This one took 4 tens to get to an integer, so it would be represented as a fraction as: 76593/(10*10*10*10)=76593/10000

OpenStudy (thecuteone25):

ok so if I do the same thing to 6.52 then I will get the answer

OpenStudy (anonymous):

Yup.

OpenStudy (thecuteone25):

so I do 6.52*10

OpenStudy (thecuteone25):

thanks I get it

OpenStudy (anonymous):

You're welcome :)